\begin{thebibliography}{9}
\bibitem{Suz} M. Suzuki, Phys.\ Lett.\ A {\bf 146} (1990) 319.
\bibitem{Yos} H. Yoshida, Phys.\ Lett.\ A {\bf 150} (1990) 262.
\bibitem{Con} M. Suzuki, Commun.\ Math.\ Phys.\ {\bf 163} (1994) 491.
\bibitem{epl} The function $F(x)$ can be given in a compact form as  
$ 2 \arctan (x / 2 \sqrt{1-x^2/4} )$ $/ x \sqrt{1-x^2/4} $ for $0<|x|<2$.
\bibitem{Gld} K. Goldberg, Duke Math.\ J. {\bf 23} (1956) 13.
\bibitem{Reu} C. Reutenauer, {\it Free Lie Algebras} (Oxford, New York, 1993),
Chap.\ 3.
\bibitem{KHS} H. Kobayashi, N. Hatano, and M. Suzuki, Physica A 
{\bf 250} (1998) 535.
\end{thebibliography}

\documentstyle[12pt]{article}
\title{Anomaly in Symplectic Integrator}
\author{Hiroto Kobayashi \\ \\ 
Department of Natural Science and Mathematics, \\
Chubu University, Kasugai 487-8501, Japan}
\date{}
\begin{document}
\maketitle
\begin{abstract}
Effective Liouville operators of the first- and the second-order
symplectic integrators are obtained for the one-dimensional
harmonic-oscillator system. The operators are defined only when the
time step is less than two. Absolute values of the coordinate and the
momentum monotonically increase for large time steps.\\
PACS numbers: 05.10.-a, 02.10.Hh \\
Keywords: exponential operator, Goldberg's theorem, convergence radius, 
conserved quantity
\end{abstract}

Symplectic integrators are widely used not only for the Hamiltonian
dynamics. Higher-order ones are also studied as the product of
exponential operators \cite{Suz,Yos} and are performed in various
areas. However, the upper limit of the time step where the scheme
works is not studied yet. Suzuki \cite{Con} studied the convergence of
decompositions of exponential operators in a Banach space. Bourbaki
treated only for the complete and normed Lie algebra. In the present
Letter, we discuss the convergence of the scheme for a specific
unbounded system. 

Let us consider the following first-order symplectic integrator for
the one-dimensional harmonic oscillator described by the Hamiltonian 
${\cal H}= (p^2 + q^2)/2 $:
\begin{equation}
\left\{
\begin{array}{l}
p' = p - x q , \\
q' = q + x p'. \end{array} \right. \label{eq:SI1}
\end{equation}
No anomalies seem to exist because the above discretization scheme is
defined for any time step $x$. However, the absolute values of $p$ and
$q$ monotonically increase for $x \ge 2$.

The above anomaly is explained by the divergence of the ''effective''
Liouville operator which corresponds to the symplectic integrator. The
discretization scheme Eq.~(\ref{eq:SI1}) is expressed by
\begin{displaymath}
\left( \begin{array}{c}
p' \\
q'     \end{array} \right) =
(1+xA)(1+xB) 
\left( \begin{array}{c}
p \\
q      \end{array} \right) =
\exp(xA) \exp(xB) 
\left( \begin{array}{c}
p \\
q      \end{array} \right) 
\end{displaymath}
with two matrices 
\begin{displaymath}
A = \left(\begin{array}{rr}
0 & 0 \\
1 & 0     \end{array} \right) , \quad
B = \left(\begin{array}{rr}
0 & -1 \\
0 &  0    \end{array} \right) 
\end{displaymath}
because of $A^2=0, \  B^2=0$. The equations 
$ABA=-A$ and $BAB=-B$ hold, and then we have
\begin{equation} 
\exp(xA) \exp(xB) = 
\exp \left\{ 
x F(x) \left(A + B + \frac{x}{\, 2 \,} [A,B] \right) 
\right\}, \label{eq:res1}
\end{equation}
where the function $F(x)$ is defined by
\begin{displaymath}
F(x) = \sum_{n=0}^\infty \frac{(n!)^2}{(2n+1)!} \,  x^{2n}
\end{displaymath}
with the convergence radius two \cite{epl}.

Defining an operator $L_1$ by
\begin{displaymath}
L_1 = A + B + \frac{x}{\, 2 \,} [A,B] 
= \left(\begin{array}{rr}
\frac{x}{\, 2 \,} &          -1 \\
       1          & - \frac{x}{\, 2 \,}    \end{array} \right) ,
\end{displaymath}
we have $i F(x) L_1$ as an effective Liouville operator which
corresponds to the symplectic integrator Eq.~(\ref{eq:SI1}). When 
a conserved quantity $E_1$ is expressed as a quadratic form of $p$ and
$q$ using a real symmetry matrix $M_1$ such as
\begin{displaymath}
E_1 = 
(p \  q) M_1
\left( \begin{array}{c}
p \\
q      \end{array} \right) ,
\end{displaymath}
the matrix $M_1 L_1$ is alternate, because the time derivative of $E_1$
is obtained from
\begin{eqnarray}
\frac{d E_1}{dt} & = & 
(p \  q) {}^t L_1 F(x) M_1
\left( \begin{array}{c}
p \\
q      \end{array} \right)  +
(p \  q) M_1 F(x) L_1
\left( \begin{array}{c}
p \\
q      \end{array} \right) \nonumber \\ & = & F(x) 
(p \  q) ({}^t (M_1 L_1) + M_1 L_1)
\left( \begin{array}{c}
p \\
q      \end{array} \right). \nonumber
\end{eqnarray}
Therefore, with
\begin{displaymath}
M_1 L_1 = \frac{1}{\, 2 \,} \det(L_1) 
\left(\begin{array}{rr}
0 & -1 \\
1 &  0    \end{array} \right) ,
\end{displaymath}
we have 
\begin{displaymath}
M_1 = \frac{1}{\, 2 \,}
\left(\begin{array}{rr}
0 & -1 \\
1 &  0    \end{array} \right) 
\left(\begin{array}{rr}
- \frac{x}{\, 2 \,} &        1 \\
         -1         & \frac{x}{\, 2 \,}    \end{array} \right) 
= \frac{1}{\, 2 \,}
\left(\begin{array}{rr}
          1         & - \frac{x}{\, 2 \,}  \\
- \frac{x}{\, 2 \,} &           1          \end{array} \right) ,
\end{displaymath}
that is, 
\begin{displaymath}
E_1 = \frac{1}{\, 2 \,} (p^2 - x p q + q^2).
\end{displaymath}

Within the convergence radius, the above Liouville operator 
$i F(x) L_1$ describes continuous dynamics along an ellipse in the 
$(p,q)$-plane with a constant value of $E_1$. Divergence of $F(x)$ can
be regarded as a break down of the continuous movement. 

Here we derive Eq.~(\ref{eq:res1}).

Expanding $\log(\exp(xA) \exp(xB))$ as the formal power series of 
$A$ and $B$, the following four kinds of terms appear:
\begin{eqnarray}
\underbrace{A B A \cdots B A}_{2n+1}, & & \nonumber \\
\underbrace{B A B \cdots A B}_{2n+1}, & & \nonumber \\
\underbrace{A B A \cdots A B}_{2n+2}, & & \nonumber \\
\underbrace{B A B \cdots B A}_{2n+2}  & & \nonumber
\end{eqnarray}
because of $A^2=0, \  B^2=0$. Coefficients of the above terms are
given by Goldberg's theorem \cite{Gld} as follows:
\begin{eqnarray}
 (-1)^n \frac{(n!)^2}{(2n+1)!}   \, x^{2n+1} ,&   & \nonumber \\
 (-1)^n \frac{(n!)^2}{(2n+1)!}   \, x^{2n+1} ,&   & \nonumber \\
 (-1)^n \frac{n!(n+1)!}{(2n+2)!} \, x^{2n+2} & = & 
 (-1)^n \frac{(n!)^2}{2(2n+1)!}  \, x^{2n+2} , \nonumber \\
-(-1)^n \frac{n!(n+1)!}{(2n+2)!} \, x^{2n+2} & = &
-(-1)^n \frac{(n!)^2}{2(2n+1)!}  \, x^{2n+2} . \nonumber
\end{eqnarray}
Because of $ABA=-A, \  BAB=-B$, we obtain 
\begin{eqnarray}
\underbrace{A B A \cdots B A}_{2n+1} & = & (-1)^n A  , \nonumber \\
\underbrace{B A B \cdots A B}_{2n+1} & = & (-1)^n B  , \nonumber \\
\underbrace{A B A \cdots A B}_{2n+2} & = & (-1)^n A B, \nonumber \\
\underbrace{B A B \cdots B A}_{2n+2} & = & (-1)^n B A, \nonumber
\end{eqnarray}
and we arrive at
\begin{eqnarray}
&   & \log(\exp(xA) \exp(xB))  \nonumber \\
& = & \sum_{n=0}^\infty \left\{
  \frac{(n!)^2}{(2n+1)!}   x^{2n+1} A + 
  \frac{(n!)^2}{(2n+1)!}   x^{2n+1} B +  
  \frac{(n!)^2}{2(2n+1)!}  x^{2n+2} A B 
- \frac{(n!)^2}{2(2n+1)!}  x^{2n+2} B A   \right\} \nonumber \\
& = & \sum_{n=0}^\infty \frac{(n!)^2}{(2n+1)!}  \,  x^{2n+1}  
\left(A + B + \frac{x}{\, 2 \,} [A,B] \right) \nonumber \\
& = & x F(x) \left(A + B + \frac{x}{\, 2 \,} [A,B] \right). \nonumber
\end{eqnarray}

Note that $\log(\exp(xA)\exp(xB))$ is expressed only with $A$, $B$,
and $[A,B]$ because of 
\begin{displaymath}
[A,[A,B]] = 2A, \quad [B,[A,B]] = - 2B.
\end{displaymath}

Next we consider the second-order symplectic integrator, which is
expressed with the two matrices $A$ and $B$ as
\begin{eqnarray}
\left( \begin{array}{c}
p' \\
q'     \end{array} \right) & = &
\left(1+\frac{x}{\, 2 \,}B \right) (1+xA) \left(1+\frac{x}{\, 2 \,}B \right) 
\left( \begin{array}{c}
p \\
q      \end{array} \right) \nonumber \\ & = &
\exp \left(\frac{x}{\, 2 \,}B \right) \exp(xA) \exp \left(\frac{x}{\, 2 \,}B \right) 
\left( \begin{array}{c}
p \\
q      \end{array} \right) . \label{eq:SI2}
\end{eqnarray}
Let us define $S(x)$ by
\begin{displaymath}
S(x) = \exp \left(\frac{x}{\, 2 \,}B \right) \exp(xA) 
\exp \left(\frac{x}{\, 2 \,}B \right) ,
\end{displaymath}
we obtain $S(-x) = S(x)^{-1}$ because of $S(x) S(-x) = 1$.
When we express $S(x)$ as
\begin{displaymath}
S(x) = \exp f(x),
\end{displaymath}
the following equation holds:
\begin{displaymath}
\exp f(-x) = S(-x) = S(x)^{-1} = \exp \{ -f(x) \},
\end{displaymath}
which indicates that $f(x)$ is an odd function. Therefore, when we
expand $f(x)$ as the formal power series of $A$ and $B$, the following
two kinds of terms appear:
\begin{eqnarray}
\underbrace{A B A \cdots B A}_{2n+1}, & & \label{eq:A} \\
\underbrace{B A B \cdots A B}_{2n+1}. & & \label{eq:B} 
\end{eqnarray}
Because of $ABA=-A, \  BAB=-B$, we can express $f(x)$ as
\begin{displaymath}
f(x) = x ( F_1 (x) A + F_2 (x) B )
\end{displaymath}
with two even functions $F_1 (x)$ and $F_2 (x)$. 

When we expand $\log (\exp X_1 \exp X_2 \exp X_3) $ as the formal
power series of $X_1$, $X_2$, and $X_3$, coefficients of the terms are
given by an extended version of Goldberg's theorem \cite{Reu,KHS}.
Because we later replace $X_1$, $X_2$, and $X_3$ as
\begin{eqnarray}
X_1 = X_3 & = & \frac{x}{\, 2 \,}B, \nonumber \\
X_2 & = & xA, \nonumber
\end{eqnarray}
two kinds of terms
\begin{eqnarray}
& & \underbrace{X_2 X_{i_1} X_2 X_{i_2} X_2 X_{i_3} X_2 
\cdots X_2 X_{i_n} X_2}_{2n+1}, \label{eq:X2} \\
& & \underbrace{X_{i_1} X_2 X_{i_2} X_2 X_{i_3} X_2 
\cdots X_2 X_{i_n} X_2 X_{i_{n+1}} }_{2n+1}  \label{eq:Xi}
\end{eqnarray}
contribute to the terms (\ref{eq:A}) and (\ref{eq:B}), respectively,
with $i_j = 1$ or $3$. According to Goldberg's theorem, coefficients
do not depend whether $X_{i_j}$ between two $X_2$'s is $X_1$ or $X_3$.
Then the coefficient of the term (\ref{eq:X2}) is given by
\begin{displaymath}
(-1)^n \frac{(n!)^2}{(2n+1)!}
\end{displaymath}
for all $X_{i_j}$. The coefficient of the term (\ref{eq:Xi}) is also
given by
\begin{eqnarray}
(-1)^n \frac{(n!)^2}{(2n+1)!} & {\rm for} & 
(i_1,i_{n+1})=(1,1), \  (3,3), \nonumber \\
(-1)^{n+1} \frac{(n-1)!(n+1)!}{(2n+1)!} & {\rm for} & 
(i_1,i_{n+1})=(1,3), \  (3,1), \nonumber
\end{eqnarray}
if $n \ge 1$. Replacing $X_1$, $X_2$, and $X_3$ with $A$ and $B$, and
using $ABA=-A$ and $BAB=-B$, we arrive at
\begin{eqnarray}
F_1 (x) & = & \sum_{n=0}^\infty \frac{(n!)^2}{(2n+1)!} \, x^{2n} =
F(x), \nonumber \\
F_2 (x) & = & 1 - \sum_{n=1}^\infty \frac{(n-1)!n!}{2(2n+1)!} \, x^{2n}
          = \left( 1 - \frac{x^2}{\,  4  \, } \right) F(x), \nonumber
\end{eqnarray}
that is,
\begin{displaymath}
\exp \left(\frac{x}{\, 2 \,}B \right) \exp(xA) 
\exp \left(\frac{x}{\, 2 \,}B \right) =
\exp \left\{ x F(x) \left(
A + \left( 1 - \frac{x^2}{\,  4  \, } \right) B \right) \right\} .
\end{displaymath}

As for the first-order scheme, defining $ L_2 $ by
\begin{displaymath}
L_2 = A + \left( 1 - \frac{x^2}{\,  4  \, } \right) B,
\end{displaymath}
we can regard $i F(x) L_2$ as an effective Liouville operator which
corresponds to the second-order symplectic integrator
Eq.~(\ref{eq:SI2}). Then we obtain the conserved quantity $E_2$ as
\begin{displaymath}
E_2 = \frac{1}{\, 2 \,} \left( p^2 + 
               \left( 1 - \frac{x^2}{\,  4  \, } \right) q^2 \right).
\end{displaymath}

Finally we note that the exponential function absolutely converges and
the product of exponentials always exists, which corresponds to the
fact that the discretization scheme as Eqs.~(\ref{eq:SI1}) and
(\ref{eq:SI2}) can be defined for all time steps. However, the
absolute values of $p$ and $q$ monotonically increase for large time
steps although Jacobian of the time-evolution operator is unity. 

I thank Profs.\ M. Suzuki and N. Hatano for their useful comments and
suggestions. I am also grateful to Dr.\ H. Watanabe for fruitful
discussions.

\end{document}